# Defect design in ferroelectrics – new insights on agglomeration

### v3

@@ 4 Conclusion
While it is commonly accepted that point defects and defect complexes in ferroelectric perovskites modify ferroelectric phase diagrams and field hysteresis, the impact of defect distribution and agglomeration is so far underrepresented. In this study, we have used $ab\ initio$ based simulation to close this knowledge gap.
@@ Appendix A Impact of Sr-substitution
In the following, we compare the impact of defects on BaTiO3 with and without Sr substitution. Analogous to the results for (Ba,Sr)TiO3 presented in Fig. 4 and Fig. 6, Figure 10 shows for BaTiO3 that non-polar defects randomly distributed in 3D mainly reduce $T_{C}^{\text{C-T}}$ (45 K, with 5 K resolution, for 2%), while their distribution in planes reduces $T_{C}^{\text{O-R}}$ (100 K for 2%) and the thermal hysteresis of the transition. Polar defects ($u_{D}=0.1$ $\hat{z}$) randomly distributed in 3D increase $T_{C}^{\text{C-T}}$ (by 45 K for 2%) and reduce $T_{C}^{\text{T-O}}$ and $T_{C}^{\text{O-R}}$ by 55 K and 25 K in cooling simulations, and by 85 K or 70 K in heating simulation, respectively. Polar defects randomly distributed in 2D, increase $T_{C}^{\text{T-O}}$ by 5 K and 40 K and decreases $T_{C}^{\text{O-R}}$ by 10 K and 15 K, during heating and cooling, respectively.
@@
Figure 10: Temperature - defect concentration phase diagrams of BaTiO3 in the presence of (a) non-polar defect dipoles ($|u_{D}|=$0.0 Å) and (b) polar defects ($|u_{D}|=$0.1 Å$\hat{x}$) for randomly distributed defects in 2D (solid lines) and in 3D (dashed lines). Red and blue lines show transition temperatures for heating and cooling, respectively.
@@
Figure 11: Changes of the field hysteresis for BaTiO3 at $T=350$ K if going from the pristine material (red) to systems with different distribution of 1 % defects (green: randomly in $\hat{z}$-planes; blue: randomly) for (a) non-polar defects and (b)–(c) polar defect dipoles ($|u_{D}|=$0.1 Å$\hat{z}$) for the electric field (b) perpendicular ($E_{x}$) and (c) collinear ($E_{z}$) to the defect dipoles. Only the polarization component $P_{i}$ along the field direction is shown.
@@
Figure 13: Comparisons of (a) field hysteresis and (b) the energy density which can be reversibly stored ($W_{\text{rec}}$) between pristine Ba0.7Sr0.3O3 in its ferroelectric ($T=250$ K, red) and paraelectric ($T=350$ K, cyan) phases to Ba0.7Sr0.3O3 with 1 % defects ($|u_{D}|=$0.1 Å$\hat{z}$) at 250 K for defect stripes (green) and random distribution in 2D (blue) in $\hat{z}$-planes. Only the polarization component $P_{x}$ along the field direction is shown in (a).
The energy which can be reversibly stored in the material depends on the full field-hysteresis rather than on this small field response, see Eqns. (2)–(3). Figure 13 compares the field hystereses and $W_{\text{rec}}$ of Ba0.7Sr0.3TiO3 between the ferroelectric ($T=250$ K, red) and paraelectric ($T=350$ K, cyan) phases of the pristine material with those in the presence of 1% polar defects ($|u_{D}|=$0.1 Å$\hat{z}$ in $z$-planes) in the ferroelectric phase (green: stripe, blue: randomly in 2D).
@@ Appendix B Relative direction of 2D ordering and polarization
Figure 14: Temperature - defect concentration phase diagrams in the presence of (a) polar defect dipoles ($|u_{D}|=$0.1 Å$\hat{x}$) and (b) non-polar defects ($|u_{D}|=$0.0 Å) for stripes of defects (solid lines) and random distribution in 2D (dashed lines). Red and blue lines show transition temperatures for heating and cooling, respectively. Numbers give the approximate number of active surface units in R phases. Note that the defect agglomerates have no effective active surface area for T and O phases as discussed in the text. Subfigure (b) is the same as shown in Fig. 4 in the main paper.
Figure 14 compares phase diagrams in the presence of (a) polar defect dipoles in the plane ($|u_{D}|=0.1$$\hat{x}$) and (b) non-polar defects ($|u_{D}|=0$) for 2D agglomerates (solid lines) and random distribution in 2D (dashed lines). While both defects have similar active surface areas and thus induce similar trend in $T_{C}^{\text{O-R}}$, the in-plane polar defects only increase $T_{C}^{\text{C-T}}$ by 11 K, decrease $T_{C}^{\text{T-O}}$ by 19 K (for 2% defects during heating), and reduce thermal hysteresis for T-O transition due to stronger internal bias fields.
@@ Appendix C Defect configurations
Figure 15: Changes of total energy induced by different configurations of 1% for non-polar ($|u_{D}|=$0.0 Å, blue) and polar ($|u_{D}|=$0.1 Å$\hat{z}$, red) defects in the (a) R phase ($T=150$ K), (b) O phase ($200$ K), and (c) T phase ($250$ K). Mean values of 20000 time steps and their standard deviations are given.
@@
Figure 16: Changes of polarization components ($|{\left\langle{P_{i}}\right\rangle}|$, where $i:x,y,z$) in cooling simulations induced by (a) non-polar defects ($|u_{D}|=$0.0 Å) and (b) polar defects ($|u_{D}|=$0.1 Å$\hat{z}$) concentrated in one plane (red) or distributed over 2 layers (blue).
@@
Figure 17: Changes of polarization components ($|{\left\langle{P_{i}}\right\rangle}|$, where $i:x,y,z$) in cooling simulations induced by (a)/(c) non-polar defects ($|u_{D}|=$0.0 Å) and (b)/(d) polar defects ($|u_{D}|=$0.1 Å$\hat{z}$) in 2D stripes for (a)–(b) 1 stripe of $w=6$ u.c. (red), 2 stripes of $w=6$ u.c. not separated ($\Delta z=0$, blue) or with a gap ($\Delta z=1$, green), and (c)–(d) 2 stripes of $w=6$ u.c. with $\Delta x=0$ (red) or $\Delta x=1$ (blue).
@@ Appendix D Defect-induced changes of polarization and strain
Figure 22: Impact of the strength of 1 % defects ($|u_{D}|$ $\hat{z}$) on the changes of polarization components $|{\left\langle{P_{i}}\right\rangle}|$, with $i:x,y,z$, in cooling simulations for (a) random distribution in 2D and (b) stripes. Colors encode the strength of the defect dipoles from 0.00 (red), 0.02 (blue), 0.03 (green), 0.04 (light blue), 0.06 (light green) to 0.10 Å (purple). (c)–(d) Histograms of local polarization at 300 K for $|u_{D}|=0.1$ $\hat{z}$ for (c) random distribution in 2D and (d) stripes.
The impact of the dipole strength on the macroscopic polarization is summarized in Fig. 22 for 1 % defects (a) randomly distributed in 2D and (b) in stripes of $w=24$ u.c. Subfigures (c) and (d) show the corresponding histograms of the local polarization for the example of $|u_{D}|=$0.1 Å at 300 K. In case of random distribution, the system is in a single domain T phase, with one peak of $P_{z}$ centered around 20 $\mu C/cm^{2}$. In case of defect stripes, this peak shows broadened distribution due to polarization rotation around the stripes. Both cases show a second small peak at 25.6 $\mu C/cm^{2}$, which belongs to the defect dipoles. The cross-over as shown in Fig. 7 (b) is therefore determined by the relative strength of the free and defect dipoles. For weak defects ($|u_{D}|\leq 0.03$$\hat{z}$), the defect dipole strength is smaller than the free dipoles, thus depolarization fields dominate. Once the strength of the defect dipole is larger than the typical values of free dipoles at slightly above $T_{C}^{\text{T-C}}$, these defects start to act as internal bias fields.
@@ Appendix E Field hysteresis
Figure 24: Field dependence of polarization component $P_{i}$ with $i=x,y,z$ at $T=250$ K with 1 % defects for (a,c) stripes and (b,d) randomly in one plane for (a,b) non-polar defects and (c,d) polar defects with ($|u_{D}|=$0.1 Å$\hat{z}$) for the electric field (a,b) along the $z$ or (c,d) perpendicular ($E_{x}$) to the defect dipoles.
@@
Figure 25: Snapshots of dipoles at 250 K for 1 % defects ($|u_{D}|=$0.1 Å$\hat{z}$) in (a) cubic agglomerate of $a=10$ u.c., (b) 3D random, (c) 2D agglomerate of $w=24$ u.c. and (d) 2D random in a plane with the applied field $-150$ kV/cm $\hat{z}$. Colors indicate the sign of $P_{z}$s. Note that all these snapshots are not time-averaged and are thus noisy.

### v4

@@ 4 Conclusion
While it is commonly accepted that point defects and defect complexes in ferroelectric perovskites modify ferroelectric phase diagrams and field hysteresis, the impact of defect distribution and agglomeration is so far underrepresented. In this study, we have used abinitioab\ initio based simulation to close this knowledge gap.
@@ Appendix A Impact of Sr-substitution
In the following, we compare the impact of defects on BaTiO3 with and without Sr substitution. Analogous to the results for (Ba,Sr)TiO3 presented in Fig. 4 and Fig. 6, Figure 10 shows for BaTiO3 that non-polar defects randomly distributed in 3D mainly reduce $T_{C}^{\text{C-T}}$ (45 K, with 5 K resolution, for 2%), while their distribution in planes reduces $T_{C}^{\text{O-R}}$ (100 K for 2%) and the thermal hysteresis of the transition. Polar defects ($u_{D}=0.1$ Å $\hat{z}$) randomly distributed in 3D increase $T_{C}^{\text{C-T}}$ (by 45 K for 2%) and reduce $T_{C}^{\text{T-O}}$ and $T_{C}^{\text{O-R}}$ by 55 K and 25 K in cooling simulations, and by 85 K or 70 K in heating simulation, respectively. Polar defects randomly distributed in 2D, increase $T_{C}^{\text{T-O}}$ by 5 K and 40 K and decreases $T_{C}^{\text{O-R}}$ by 10 K and 15 K, during heating and cooling, respectively.
@@
Figure 10: Temperature - defect concentration phase diagrams of BaTiO3 in the presence of (a) non-polar defect dipoles ($|u_{D}|=$0.0 Å) and (b) polar defects ($|u_{D}|=$0.1 $\hat{x}$) for randomly distributed defects in 2D (solid lines) and in 3D (dashed lines). Red and blue lines show transition temperatures for heating and cooling, respectively.
@@
Figure 11: Changes of the field hysteresis for BaTiO3 at $T=350$ K if going from the pristine material (red) to systems with different distribution of 1 % defects (green: randomly in $\hat{z}-$planes; blue: randomly) for (a) non-polar defects and (b)–(c) polar defect dipoles ($|u_{D}|=$0.1 $\hat{z}$) for the electric field (b) perpendicular ($E_{x}$) and (c) collinear ($E_{z}$) to the defect dipoles. Only the polarization component $P_{i}$ along the field direction is shown.
@@
Figure 13: Comparisons of (a) field hysteresis and (b) the energy density which can be reversibly stored ($W_{\text{rec}}$) between pristine Ba0.7Sr0.3O3 in its ferroelectric ($T=250$ K, red) and paraelectric ($T=350$ K, cyan) phases to Ba0.7Sr0.3O3 with 1 % defects ($|u_{D}|=$0.1 $\hat{z}$) at 250 K for defect stripes (green) and random distribution in 2D (blue) in $\hat{z}-$planes. Only the polarization component $P_{x}$ along the field direction is shown in (a).
The energy which can be reversibly stored in the material depends on the full field-hysteresis rather than on this small field response, see Eqns. (2)–(3). Figure 13 compares the field hystereses and $W_{\text{rec}}$ of Ba0.7Sr0.3TiO3 between the ferroelectric ($T=250$ K, red) and paraelectric ($T=350$ K, cyan) phases of the pristine material with those in the presence of 1% polar defects ($|u_{D}|=$0.1 $\hat{z}$ in $z$-planes) in the ferroelectric phase (green: stripe, blue: randomly in 2D).
@@ Appendix B Relative direction of 2D ordering and polarization
Figure 14: Temperature - defect concentration phase diagrams in the presence of (a) polar defect dipoles ($|u_{D}|=$0.1 $\hat{x}$) and (b) non-polar defects ($|u_{D}|=$0.0 Å) for stripes of defects (solid lines) and random distribution in 2D (dashed lines). Red and blue lines show transition temperatures for heating and cooling, respectively. Numbers give the approximate number of active surface units in R phases. Note that the defect agglomerates have no effective active surface area for T and O phases as discussed in the text. Subfigure (b) is the same as shown in Fig. 4 in the main paper.
Figure 14 compares phase diagrams in the presence of (a) polar defect dipoles in the plane ($|u_{D}|=0.1$ $\hat{x}$) and (b) non-polar defects ($|u_{D}|=0$ Å) for 2D agglomerates (solid lines) and random distribution in 2D (dashed lines). While both defects have similar active surface areas and thus induce similar trend in $T_{C}^{\text{O-R}}$, the in-plane polar defects only increase $T_{C}^{\text{C-T}}$ by 11 K, decrease $T_{C}^{\text{T-O}}$ by 19 K (for 2% defects during heating), and reduce thermal hysteresis for T-O transition due to stronger internal bias fields.
@@ Appendix C Defect configurations
Figure 15: Changes of total energy induced by different configurations of 1% for non-polar ($|u_{D}|=$0.0 Å, blue) and polar ($|u_{D}|=$0.1 $\hat{z}$, red) defects in the (a) R phase ($T=150$ K), (b) O phase ($200$ K), and (c) T phase ($250$ K). Mean values of 20000 time steps and their standard deviations are given.
@@
Figure 16: Changes of polarization components ($|{\left\langle{P_{i}}\right\rangle}|$, where $i:x,y,z$) in cooling simulations induced by (a) non-polar defects ($|u_{D}|=$0.0 Å) and (b) polar defects ($|u_{D}|=$0.1 $\hat{z}$) concentrated in one plane (red) or distributed over 2 layers (blue).
@@
Figure 17: Changes of polarization components ($|{\left\langle{P_{i}}\right\rangle}|$, where $i:x,y,z$) in cooling simulations induced by (a)/(c) non-polar defects ($|u_{D}|=$0.0 Å) and (b)/(d) polar defects ($|u_{D}|=$0.1 $\hat{z}$) in 2D stripes for (a)–(b) 1 stripe of $w=6$ u.c. (red), 2 stripes of $w=6$ u.c. not separated ($\Delta z=0$, blue) or with a gap ($\Delta z=1$, green), and (c)–(d) 2 stripes of $w=6$ u.c. with $\Delta x=0$ (red) or $\Delta x=1$ (blue).
@@ Appendix D Defect-induced changes of polarization and strain
Figure 22: Impact of the strength of 1 % defects ($|u_{D}|$ $\hat{z}$) on the changes of polarization components $|{\left\langle{P_{i}}\right\rangle}|$, with $i:x,y,z$, in cooling simulations for (a) random distribution in 2D and (b) stripes. Colors encode the strength of the defect dipoles from 0.00 (red), 0.02 (blue), 0.03 (green), 0.04 (light blue), 0.06 (light green) to 0.10 Å (purple). (c)–(d) Histograms of local polarization at 300 K for $|u_{D}|=0.1$ Å $\hat{z}$ for (c) random distribution in 2D and (d) stripes.
[…]
@@ Appendix E Field hysteresis
 $E_{x}$ (kV/cm). $P_{z}$$P_{y}$$P_{x}$(d)
Figure 24: Field dependence of polarization component $P_{i}$ with $i=x,y,z$ at $T=250$ K with 1 % defects for (a,c) stripes and (b,d) randomly in one plane for (a,b) non-polar defects and (c,d) polar defects with ($|u_{D}|=$0.1 $\hat{z}$) for the electric field (a,b) along the $z$ or (c,d) perpendicular ($E_{x}$) to the defect dipoles.
@@
Figure 25: Snapshots of dipoles at 250 K for 1 % defects ($|u_{D}|=$0.1 $\hat{z}$) in (a) cubic agglomerate of $a=10$ u.c., (b) 3D random, (c) 2D agglomerate of $w=24$ u.c. and (d) 2D random in a plane with the applied field 150-150 kV/cm $\hat{z}$. Colors indicate the sign of $P_{z}$s. Note that all these snapshots are not time-averaged and are thus noisy.